OpenStudy (anonymous):

Simplify the expression. (-1^5/-2^-3)^2

OpenStudy (anonymous):

your equation simplifies to (-1^5/1/-2^3)^2 so -2^3 = -8 now you have (-1^5/1/-8)^2 which is (-8*-1^5) on the top, over 1 so 8/1. Which is just 8. Then square it and you get 64

OpenStudy (anonymous):

Awesome ! Thank you. Medal given

OpenStudy (anonymous):

just joined this site, and watched this thread. Are you sure thats the answer?

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 2 } \right)^{16}\]

OpenStudy (mathstudent55):

\( \left( \dfrac{-1^5}{-2^{-3}} \right)^2 \) \( =\left( \dfrac{-1}{-\frac{1}{8}} \right)^2 \) \( = 8^2 = 64\)

OpenStudy (anonymous):

just revised now,, can also be done as : \[\left( \frac{ 2^{3} }{ 1 }\ \times \frac{ 1 }{ 1^{5} } \right) ^{2} = 8^{2} = 64\]